# Asymmetry Analysis of Bilateral Shapes

### v1

@@ Abstract
Many biological objects possess bilateral symmetry about a midline or midplane, up to a “noise” term. This paper uses landmark-based methods to measure departures from bilateral symmetry, especially for the two-group problem where one group is more asymmetric than the other. In this paper, we formulate our work in the framework of size-and-shape analysis including registration via rigid body motion. Our starting point is a vector of elementary asymmetry features defined at the individual landmark coordinates for each object. We introduce two approaches for testing. In the first, the elementary features are combined into a scalar composite asymmetry measure for each object. Then standard univariate tests can be used to compare the two groups. In the second approach, a univariate test statistic is constructed for each elementary feature. The maximum of these statistics lead to an overall test statistic to compare the two groups and we then provide a technique to extract the important features from the landmark data. Our methodology is illustrated on a pre-registered smile dataset collected to assess the success of cleft lip surgery on human subjects. The asymmetry in a group of cleft lip subjects is compared to a group of normal subjects, and statistically significant differences have been found by univariate tests in the first approach. Further, our feature extraction method leads to an anatomically plausible set of landmarks for medical applications.
Keywords— asymmetry scores, cleft lip disfigurement, facial trauma surgery, feature selection, registration, shape analysis, size-and-shape, smile.
@@ 1 Introduction
Bilateral symmetry (left-right symmetry) is a key property in many biological settings. An object (treated as a set of unlabelled points known as landmarks) is said to be bilaterally symmetric if its reflection through some midline or midplane $\mathcal{P}$ is exactly the same as the original object. In practice, exact bilateral symmetry seldom holds and it is of interest to study the extent of any asymmetry. Research about the symmetries of objects in the real world has a long history (see, for example, Weyl, 1952). Of particular interest is the hypothesis of bilateral symmetry, especially in biology (Palmer and Strobeck, 1986). Statistical methods to test asymmetry were first formalized in Mardia et al. (2000). Subsequently, a wide variety of statistical methods have been developed; see, for examples, Kent and Mardia (2001), Bock and Bowman (2006), Ajmera et al. (2022) and Ajmera et al. (2023).
@@
The focus of this paper is on landmark-based objects as in Mardia et al. (2000). Each object is a set of $K$ labelled points or landmarks, represented as a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions. The cases $M=2$ and $M=3$ are the most important in practice. Call an object bilateral if its landmarks can be divided into two categories: pairs and solos, where paired landmarks locate on both sides of $\mathcal{P}$, while solos are unpaired. For a bilaterally symmetric object the two landmarks in a pair are equally spaced about the midplane and the solos lie on the midplane. If there are $K_{P}$ pairs of landmarks and $K_{S}$ solo landmarks, then the total number of landmarks is $K=2K_{P}+K_{S}$. If the dataset has not been pre-registered, then registration is required. The details of registration are given in Section 2.1.
@@
The asymmetry information in an object can be represented as a vector of coordinatewise elementary asymmetry features, $\bm{d}$, say, defined in more details in Section 2.2. Increasing the magnitude of any element of this vector indicates greater asymmetry. The elementary feature vector is also the starting point of Bock and Bowman (2006) and Patel et al. (2023) but we define its elements explicitly in Section 2.2.
Given two groups of objects, a natural question is whether one group is more asymmetric than the other. Section 3 looks at two testing strategies. In the first strategy the elements of the feature vector $\bm{d}$ are combined into a single number called a composite score. Then a univariate test such as the two-sample t-test or the Mann-Whitney U test can be used to compare the two groups. The use of a univariate summary statistic for each object facilitates the comparison between objects and between groups of objects. The mathematical details are given in Section 3.1, our composite score extends the asymmetry measures used by Bock and Bowman (2006) and Patel et al. (2023).
In the second strategy, the two groups are compared using each of the elements in $\bm{d}$, yielding a collection of test statistics. The maximum of these test statistics gives an overall test statistic that can be used to test for a difference between the two groups. This approach is similar to the union-intersection test (UIT) in multivariate analysis, see, for example, Mardia et al. (2024). In particular, when the null hypothesis is rejected, it is possible to investigate which elementary features are responsible. The mathematical details are given in Section 3.2.
The motivation for the paper is given by a set of smile data on human subjects. Each subject was asked to smile, and a four-dimensional movie was made of the face (three spatial dimensions plus time using Di4D motion capture system). A set of landmarks was identified around the lips and followed through the frames of the movie. The subjects come from two groups: normal subjects and subjects who have had cleft lip surgery. The objective is to assess how successful the surgery has been; that is, are the cleft lip subjects any more asymmetric than normal people. The details of the smile data are provided in Section 4. The data is pre-registered and the objects at three fixed frames (the beginning, middle and end of the smile) are analyzed in Section 4 following Patel et al. (2023). We end the paper with some discussion in Section 5. The procedures for estimating the midline or midplane are given in Appendix A.
@@ 2.1 Registration
We will introduce the following notations: let $X$, with elements $X[k,m]$, be a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions for a single object or subject or individual. We write as $X[k,\ ]\in\mathbb{R}^{M}$ for the coordinates of $k$th landmark, $k=1,\dots,K$. The most important choice for $M$ is $M=3$, and this case is emphasized in the presentation, but the mathematical theory is valid for any $M\geq 1$. In general, upper case letters are used for matrices and bold lower case letters for vectors. In addition the letters $K,M,J$ are reserved for the number of landmarks, number of dimensions, and number of elementary features (Section 2.2), respectively, with typical indices given by the lower case letters $k,m,j$. In general, each data object $X$ will be viewed as a noisy version of a bilaterally symmetric configuration.
Rigid body transformations If $X$ is a configuration, then a rigid body motion takes $X$ to $\bm{1}_{M}\bm{c}^{T}+XR=X^{*}$, say, where $\bm{c}$ is an $M$-dimensional translation vector and $R$ is an $M\times M$ rotation matrix. For the purposes of this paper, we say that $X$ and $X^{*}$ have the same size-and-shape following the standard terminology in shape analysis (see, for example, Dryden and Mardia (2016)). That is, the size-and-shape of $X$ is the equivalence class of configurations under the group of rigid body motions. The asymmetry information carried by the size-and-shape is of interest and such information is unaffected under rigid body motion. The size is emphasized here, since the size information is taken into consideration. Recall that the term “shape” is a standard terminology in statistical shape analysis which means the equivalence class under the larger group of similarity transformations so the scale is also filtered out.
@@
Two types of Registration For every bilateral data object $X$, suppose there is an associated midplane $\mathcal{P}$. Ways in which $\mathcal{P}$ might be determined are disussed below. For data analysis, it is helpful to “register” $X$ using a rigid body motion so that after transformation $\bm{n}=\bm{e}_{1}$ and $b=0$, i.e., the midplane passes through the origin and the normal direction is given by the first coordinate axis, while still keeps the shape of $X$ unchanged.
We will distinguish two types of registration:
  1. 1.
    Axis registration
  2. 2.
@@
Axis registration takes the form described in the previous paragraph. In terms of the human head, after registration, the left-right axis lies along the first coordinate axis. However, the position of the down-up and back-front axes are not clearly distinguished from one another. It is possible to tilt the head forwards or backwards while still preserving axis registration.
In some settings it is possible to carry out a stronger version of registration, we will call basis registration, in which all the coordinate axes have a natural interpretation. For example, for 3D measurements of the human head, the midplane is known as the sagittal plane, and it is natural to require the coordinate axes to have the following interpretations:
  • coordinate 1: left-right
  • coordinate 2: down-up
  • coordinate 3: back-front
Figure 1 shows these coordinate axes on a human face. Further details are given in Section 4.
Figure 1: Coordinates system for human face with principal planes. (Source: Patel et al. (2023); coordinate labels have been corrected.)
To carry out axis registration it is necessary to know or estimate the midplane. There are three typical scenarios: (a) the midplane is known a priori; (b) the midplane has been estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) the midplane is estimated from the landmark positions using Procrustes analysis. Scenario (c) is described in detail in Appendix A following Mardia et al. (2000) and Kent and Mardia (2001).
To carry out basis registration, it is also necessary to know or estimate the $M-1$ remaining meaningful basis vectors in $\mathcal{P}$. In this case the three possibilities become: (a) the basis is known a priori; (b) the basis is estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) hybrid Procrustes analysis. For details in Procrustes analysis, see, for example Dryden and Mardia (2016). By “hybrid”, we mean that (i) the normal direction to the midplane is estimated by combining an object with its reflection and using generalized Procrustes analysis (GPA), (ii) an “average shape” is found from all the objects using GPA, (iii) the midplane for the average shape is oriented so the basis directions have natural interpretations using expert knowledge, and (iv) each data configuration is aligned to the average shape using ordinary Procrustes analysis. The details are given in Appendix A.
For the smile data application in Section 4, expert knowledge (scenario (b)) has been used to estimate the midplane with the stated interpretations for the coordinate axes as above in basis registration.
@@ 2.2 Elementary features
Suppose the size-and-shape dataset has been basis registered about a midplane $\mathcal{P}$ that passes through the origin and is normal to the first coordinate axis. Then it is possible to define various elementary features to describe the asymmetry of each data object. Paired and solo landmarks are treated separately.
Let $K_{L}$ and $K_{R}$ denote the total number of landmarks on the left or right of $\mathcal{P}$. We have $K_{L}=K_{R}=K_{P}$. Let $k_{L}$ and $k_{R}$ denote the left and right indices for a typical landmark pair for a bilateral object $X$ and consider the $M$ coordinatewise signed elementary features for the landmark pairs,
@@
That is, there is one feature for each coordinate and each landmark pair. So there are in total $MK_{P}$ coordinatewise features for all paired landmarks. For this specific $\mathcal{P}$, the first coordinates for the two landmarks in one pair have opposite signs, while other coordinates have the same sign. Further, for a symmetric object, the quantities in equations (2.1) and (2.2) should all be 0. Thus, taking the sum of the first coordinates and differences for other coordinates allow us to quantify the departure from symmetry. Similarly, if $k_{S}$ is a typical solo landmark, consider the single elementary feature
@@
There are in total $K_{S}$ elementary features for solos. Under bilateral symmetry with $\mathcal{P}$ chosen as above, the first coordinate of a solo landmark will be 0, with no restriction on other coordinates. Hence, a single feature as in (2.3) is adequate to describe the asymmetry of a solo landmark.
The coordinatewise elementary features can be collected into a signed elementary feature vector $\bm{d}=(d_{j})$, say, of length $J_{\text{basis}}=MK_{P}+K_{S}$. The elements of $\bm{d}$ will either be listed sequentially with a subscript index, i.e., $d_{j},\ j=1,\ldots,J_{\text{basis}}$, or using parentheses within square brackets, as in (2.1)-(2.3).
When only axis registration is available, the coordinate information in (2.2) does not have a well-defined interpretation. This is because the coordinates except the first one of paired landmarks would be changed after rotation within $\mathcal{P}$, hence leads to changes in values of (2.2) even though the asymmetry information does not change. Note that the equation (2.1) will not change. Further, at each landmark the coordinate information can be collected into a single number, called a landmark elementary feature which is defined per landmark,
$d^{*}[(k_{L},k_{R})]=\left\{\sum_{m=1}^{M}d^{2}[(k_{L},k_{R}),m]\right\}^{1/2},$ (2.4)
which does have a well-defined interpretation as the Euclidean distance between one landmark in a pair and the reflection of the other landmark in the pair. Rotation within $\mathcal{P}$ would not change this Euclidean distance. For a solo landmark, (2.3) is always well defined since the first coordinate of a solo remains unchanged after rotation within $\mathcal{P}$; hence it is both a coordinatewise and a landmark elementary feature. Thus, overall in the case of axis registration, a smaller set of $J_{\text{axis}}=K_{P}+K_{S}$ elementary features can be collected into a vector for further analysis. On the other hand, for $M=2$, there is no difference between axis and basis registration. Thus, equations (2.1)-(2.3) are all valid for both types of registrations in this case.
Another useful way to think about the elementary features is by using reflection. Let $X$ be a basis registered configuration, and let $X^{(\text{refl})}$ denote the reflection of $X$ about $\mathcal{P}$. Computationally, this matrix is obtained by changing the sign of the first column of $X$ and interchanging the row indices for each landmark pair. Then the elementary features of $X$ can be obtained by comparing the elements of $X$ and $X^{(\text{refl})}$. If $X$ is bilaterally symmetric about $\mathcal{P}$, then $X=X^{(\text{refl})}$ and elementary features vanish, i.e. $\bm{d}=\bm{0}$.
In many examples, such as the smile application studied later in this paper in Section 4, the direction of asymmetry, for example, left or right asymmetry, can vary between individuals and is not thought to be interested. We will focus on the absolute values of the elementary features. Define the absolute elementary feature vector $\bm{a}$ with elements
@@
where $J=J_{\text{basis}}$ or $J=J_{\text{axis}}$, as appropriate. In other words, only the extent of the asymmetry is interested. The elements of $\bm{a}$ can also be written in a similar way as elements of $\bm{d}$ in (2.1)-(2.3). In the rest of the paper, each configuration $X$ is reduced to an absolute elementary feature vector $\bm{a}$ for further numerical and statistical analysis.
2.3 A Simple Illustrative Example
To illustrate the construction of the vector $\bm{a}$, consider two configurations in $M=2$ dimensions with $K=4$ landmarks given by
$X_{1}=\begin{pmatrix}-1&\phantom{-}0\\ \phantom{-}0&\phantom{-}1\\ \phantom{-}1&\phantom{-}0\\ \phantom{-}0&-1\end{pmatrix},\quad X_{2}=\begin{pmatrix}-0.95&\phantom{-}0.36\\ -0.28&\phantom{-}2.11\\ \phantom{-}0.99&\phantom{-}0.54\\ -0.31&-1.37\end{pmatrix}.$ (2.6)
Suppose the landmarks consist of one landmark pair, $(k_{L},k_{R})=(1,3)$ and two solos, $k_{S}=2,4$. Hence $K_{P}=1$ and $K_{S}=2$. Further, suppose both configurations are treated as basis registered with $\mathcal{P}$ normal to the first coordinate axis, i.e. $\mathcal{P}$ is the $y$-axis.

### v2

@@ Abstract
Many biological objects possess bilateral symmetry with some noise about a midline or midplane. This paper uses landmark-based methods to measure departures from bilateral symmetry, especially for paired data and two-group problems where one group may be more symmetric than the other. We formulate our work under size-and-shape analysis framework including pre-registration under rigid body motion. We then construct a vector of elementary asymmetry features defined per landmark coordinate per object. Next, we introduce two approaches for comparing bilateral symmetry between the two groups or paired data: (1) the elementary features are combined into a scalar asymmetry measure for each object, then standard univariate tests are used to test for asymmetries; (2) a univariate test statistic is constructed for each feature, then the maximum of these statistics leads to an overall test statistic and important features can be extracted from the data. Our methodology is illustrated on two registered datasets collected to assess the success of two kinds of surgeries: (1) cleft lip surgery data, where the symmetry is compared between a group of cleft lip subjects and normal subjects; (2) orthognathic surgery data, where the patients’ symmetry after the orthognathic surgery are compared to that before surgery.
Keywords: asymmetry measures, cleft lip surgery, orthogonathic surgery, registration, shape analysis, size-and-shape analysis of smile.
@@ 1 Introduction
Bilateral symmetry, also called left-right symmetry, is a key property in many biological settings. An object is said to be bilaterally symmetric if its reflection through some midline (for objects in two dimensions) or midplane (for objects in three dimensions) passing through the center of the object is exactly the same as the original object. In practice, exact bilateral symmetry seldom holds and it is of interest to study the extent of any asymmetry. Research on the symmetries of objects in the real world has a long history (see, for example, Weyl, 1952). Statistical methods to test asymmetry were first formalized in Mardia et al. (2000). Subsequently, a wide variety of statistical methods have been developed; see, for examples, Kent and Mardia (2001), Bock and Bowman (2006), Ajmera et al. (2022) and Ajmera et al. (2023).
@@
The focus of this paper is on landmark-based objects as in Mardia et al. (2000). Each object is a set of $K$ labelled points or landmarks, represented as a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions. The cases $M=2$ and $M=3$ are the most important in practice. We call an object bilateral if its landmarks can be divided into two categories: pairs and solos. An object is bilaterally symmetric about an M1M-1-dimensional midplane $\mathcal{P}$ if all the solos lie on the midplane, and the two landmarks in each pair are equally spaced about $\mathcal{P}$, meaning that their midpoint lies on $\mathcal{P}$ and the vectors to each from their midpoint are orthogonal to $\mathcal{P}$. Let there be $K_{P}$ pairs and $K_{S}$ solos; then $K=2K_{P}+K_{S}$. After we have determined the $M$ orthogonal axes for the object, the intersection point between the midplane and the first coordinate axis can be computed. Under translation and rotation it is ensured that the midplane $\mathcal{P}$ passes through the origin and is orthogonal to the first coordinate axis. All objects in the dataset should be registered so that all of them share the same midplane $\mathcal{P}$, hence, further analyses are available. If the dataset has not been pre-registered before the analyses, then registration is required; the details are given in Section 2.1.
This paper is motivated by two medical questions related to the level of asymmetry in a person’s smile. One is related to cleft lip surgery which is performed so that form and function can be restored to particularly lip symmetry at rest and during movement. The second is related to orthognathic surgery, which is performed to correct jaw discrepancies in order to restore the form and function but the surgery can influence facial and lip symmetry.
We have been provided with appropriate data in each case. We will refer the cleft lip data as Smile Data 1, and the orthognathic surgery data as Smile Data 2. In both datasets, we have a set of landmarks on the lip area as appropriate to each case; Smile Data 1 contains 24 landmarks over the lip while Smile Data 2 contains 16 landmarks over the lip and nose. More details of the data are provided in Section 5 and Section 6 respectively. In each case, the subjects were asked to smile from rest (close lip) to maximum open lip smile and a four-dimensional (three spatial dimensions plus time) movie capture was made of the face using Di4D motion capture system (https://di4d.com). A superimposed finite element mesh combined with optical flow techniques were used to follow the landmarks through the frames of the capture, taken at one-second intervals, of the capture. Note that these include both anatomical landmarks, e.g. the corners of the mouth, and pseudo landmarks, e.g. a midpoint of the upper right lip defined at first frame and then followed through successive frames of the capture by optical flow.
Smile Data 1 contains two groups: normal subjects and subjects who had previous cleft lip surgery. Smile Data 2 includes a single group of subjects before and after orthognathic surgery, so it is a paired dataset. The aim is to assess the success of these two different surgeries, that is, are the cleft lip subjects more symmetric after the surgery, and has orthognathic surgery had an impact on symmetry. Both datasets are pre-registered in the same way and we have three fixed frames (the beginning, middle and end of the smile) for each subject.
@@
The asymmetry information in an object is represented as a vector of coordinatewise elementary asymmetry features, $\bm{d}$, say, defined in Section 2.2. Increasing the magnitude of any element in $\bm{d}$ indicates greater asymmetry. The elementary feature vector is also the starting point of Bock and Bowman (2006) and Patel et al. (2023), and we define its elements explicitly in Section 2.2.
Given two groups of objects, a natural question is whether one group is more asymmetric than the other. Section 3 looks at two testing strategies which we name combine-then-compare and compare-then-combine respectively.
Our methodology involves simultaneous comparisons, hence, meta-analysis is considered. The literature on meta-analysis as required in the paper is scattered over several papers, and in Section 4, we give an overview of meta-analysis as required for our work. The meta-analysis is carried out in Section 5 and 6 respectively for the two datasets.
The paper concludes with a discussion in Section 7.
@@ 2.1 Registration
[…]
Rigid body transformations If $X$ is a configuration, then a rigid body motion takes $X$ to $X^{*}=\bm{1}_{M}\bm{c}^{T}+XR$, say, where $\bm{c}\in\mathbb{R}^{M}$ is a translation vector and $R$ is an $M\times M$ rotation matrix. For the purposes of this paper, we say that $X$ and $X^{*}$ have the same size-and-shape, as the size-and-shape of $X$ is the equivalence class of configurations under the group of rigid body motions (see, for example, Dryden and Mardia (2016)). The asymmetry information carried by the size-and-shape is of interest and such information is unaffected under rigid body motion. The size is emphasized here, since the size information is taken into consideration. Recall that the term “shape” is a standard terminology in statistical shape analysis which means the equivalence class under the larger group of similarity transformations so the scale is also filtered out.
@@
Two types of Registration We distinguish two types of registrations
  1. (a)
    Axis registration.
  2. (b)
@@
Suppose there is an associated midplane $\mathcal{P}$ for every bilateral data object $X$. Ways in which $\mathcal{P}$ might be determined are discussed below. For data analysis, it is helpful to “register” $X$ using a rigid body motion so that after transformation $\bm{n}=\bm{e}_{1}$ and $b=0$, i.e., the midplane passes through the origin and the normal direction is given by the first coordinate axis, while still keeps the size-and-shape of $X$ unchanged.
Figure 1: Coordinates system used for registration of human face used in our illustrative smile examples: with three principal planes. In the figure at top row, the $x$-$y$ plane is in green, the $y$-$z$ plane is in red and the $x$-$z$ plane is in blue. For the three sub-figures in the bottom row, the $x$-axis is in green, $y$-axis is in red and $z$-axis is in blue.
In some settings it is possible to carry out a stronger version of registration than the axis registration, we will call basis registration, in which all the coordinate axes have a natural interpretation. In our illustrative smile examples, for 3D measurements of the human head, the midplane is known as the sagittal plane, and it is natural to require the coordinate axes to have the following interpretations:
  • (a)
    $x$-coordinate: left-right
  • (b)
    $y$-coordinate: down-up
  • (c)
    $z$-coordinate: back-front
Figure 1 shows these coordinate axes as well as the three planes on a human face which have been used in pre-registration for our smile data. That is, the $x$-axis is positioned horizontally on the face (positive direction is left from the view point of the subject), the $y$-axis is placed vertically on the face (positive direction is upward), whereas the $z$-axis is positioned in the direction inwards and outwards of the face (positive direction is outward), see Figure 1.
In general, to carry out basis registration, it is also necessary to know or estimate M1M-1 meaningful basis vectors in $\mathcal{P}$. In this case the three possibilities become: (a) the basis is known a priori; (b) the basis is estimated using expert knowledge; or (c) the basis is estimated using hybrid Procrustes analysis. For details of Procrustes analysis, see, for example, Goodall (1991) and Dryden and Mardia (2016).
For our illustrative smile examples, expert knowledge (scenario (b)) has been used to estimate the midplane with the stated interpretations for the coordinate axes as above (Figure 1). It may be noted that there will be one more reflection before the registration in some data; specifically, for those cleft lip subjects in Smile Data 1 with the cleft on the right, we reflect in the midplane so that the cleft is consistently on the left.
@@ 2.2 Elementary features
Suppose the size-and-shape dataset has been basis registered about a midplane $\mathcal{P}$ that passes through the origin and is normal to the first coordinate axis and the M1M-1 axes of $\mathcal{P}$ are all determined. Then it is possible to define various elementary features to describe the asymmetry of each data object. Paired and solo landmarks are treated separately. We begin with an idealized bilaterally symmetric object $X^{S}$. Each of the $K$ landmarks of $X^{S}$ will be either to the left of $\mathcal{P}$, to the right of $\mathcal{P}$, or on $\mathcal{P}$. The counts of landmarks in these subsets are respectively $K_{L}$, $K_{R}$, and $K_{S}$ where $K_{L}=K_{R}=K_{P}$ .
For any $K\times M$ observation $X$ modeled on $X^{S}$, landmarks, identified by their indices, follow the prescription of $X^{S}$. Let $k_{L}$ and $k_{R}$ denote the left and right indices for a typical landmark pair from a bilateral object $X$ and consider the $M$ coordinatewise signed elementary features for the landmark pairs,
@@
That is, there is one feature for each coordinate of each landmark pair, so there are in total $MK_{P}$ coordinatewise features. The first coordinates for the two landmarks in one pair have opposite signs, while other coordinates have the same sign with respect to $\mathcal{P}$. Note that this pattern is true for both axis and basis registrations. For $X^{S}$, the quantities in equations (2.1) and (2.2) should all be 0. Thus, the quantities given in (2.1) and (2.2) quantify the departure from symmetry. Similarly, if $k_{S}$ is a typical solo landmark, consider the single elementary feature
@@
There are in total $K_{S}$ elementary features for solos. For $X^{S}$, the first coordinate of a solo landmark will be 0, with no restriction on other coordinates. Hence, a single feature as in (2.3) is adequate to describe the asymmetry of a solo landmark.
The coordinatewise elementary features can be collected into a signed elementary feature vector $\bm{d}=(d_{j})$, say, of length $J_{\text{basis}}=MK_{P}+K_{S}$. The elements of $\bm{d}$ will either be listed sequentially with a subscript index, i.e., $d_{j},\ j=1,\ldots,J_{\text{basis}}$, or using parentheses within square brackets, as in (2.1)-(2.3). Note that the vector $\bm{d}$ is a function of matrix $X$ and $\mathcal{P}$.
When only axis registration is available, the coordinate information in (2.2) does not have a well-defined interpretation. This is because the coordinates except the first one of paired landmarks would be changed after rotation within $\mathcal{P}$, hence leads to changes in values of (2.2) even though the asymmetry information does not change (but note that equation (2.1) does not change). Further, at each landmark the coordinate information can be collected into a single number, called a landmark elementary feature which is defined per landmark,
$d^{*}[(k_{L},k_{R})]=\left\{\sum_{m=1}^{M}d^{2}[(k_{L},k_{R}),m]\right\}^{1/2},\ d^{*}[(k_{S})]=|X[k_{S},1]|,$ (2.4)
where $d$’s are defined by (2.1) and (2.2).
Another useful way to think about the elementary features is by using reflection. Let $X^{(\text{refl})}$ denote the reflection of $X$ about $\mathcal{P}$. Computationally, this matrix is obtained by changing the sign of the first column of $X$ and interchanging the row indices for each landmark pair. Then the elementary features of $X$ can be obtained by comparing the elements of $X$ and $X^{(\text{refl})}$. If $X$ is bilaterally symmetric about $\mathcal{P}$, then $X=X^{(\text{refl})}$ and $\bm{d}=\bm{0}$.
In many examples, such as the illustrative smile examples studied later in this paper in Section 5 and 6, the direction of asymmetry, for example, left or right asymmetry, can vary between individuals and is not thought interesting. We will focus on the absolute values of the elementary features. Define the absolute elementary feature vector $\bm{a}$ with elements
@@
where $J=J_{\text{basis}}$ or $J=J_{\text{axis}}$, as appropriate. In other words, only the extent of the asymmetry is interesting. The elements of $\bm{a}$ can also be written in a similar way as elements of $\bm{d}$ in (2.1)-(2.4). In the rest of the paper, each configuration $X$ is reduced to an absolute elementary feature vector $\bm{a}$ for further numerical and statistical analysis. Note that $\bm{a}$ is also a function of matrix $X$ and $\mathcal{P}$.
2.3 Example 1
We give here a simple example to illustrate our basic notation and construction of the vector $\bm{a}$ and related quantities. Consider two configurations in $M=2$ dimensions with $K=4$ landmarks given by